# A Phylogenetic Approach to Genomic Language Modeling

### v1

@@ 1 Introduction
In recent work, Benegas et al. [1] introduced a gLM named GPN-MSA that achieves cutting-edge performance at this task, even when compared with well-established variant effect predictors. These results are due to a unique aspect of the design of GPN-MSA: it utilizes a multiple sequence alignment (MSA) of diverse vertebrate genomes to the human reference genome as supplementary features, demonstrating that these data are an invaluable resource for modeling human genomes. However, for making a prediction, it requires as input an MSA for the same set of species as in training data, which makes it difficult to apply the model to other species or to regions where alignment to the human genome is poor. This undermines its utility for transfer learning. Furthermore, if genomes too similar to the human genome are included in the MSA, the model learns to simply copy from them, resulting in the learned probability distribution being not very useful for variant effect prediction. Therefore, most primates were excluded from GPN-MSA’s training data, which is not ideal since they provide useful information about relatively recent evolutionary constraints relevant to humans [15]. Lastly, it relies on existing conservation annotations [32, 27] to filter training data and to re-weight the loss function to mitigate the influence of poorly conserved regions on the model.
@@ 2.2 Molecular Phylogenetics
The basic data that represent evolutionary relationships between aligned sequences are phylogenetic trees. The leaves of the tree correspond to extant species that are included in the alignment. The interior nodes, on the other hand, correspond to extinct ancestral species. In our context, edges are labeled with values in $[0,\infty)$ that represent the evolutionary distances between nodes.
@@ 3.2 Loss Function
$\mathbb{P}_{\mathsf{F81}}(a\mid\theta,b,t)\geq\begin{cases}\operatorname{sigmoid}\left(\log t+\sum_{a\in\mathcal{N}}\theta_{a}\right)(\pi_{a}-1)+1&\text{if}\ a=b,\\ \operatorname{sigmoid}\left(\log t+\sum_{a\in\mathcal{N}}\theta_{a}\right)\pi_{a}&\text{otherwise},\end{cases}$ (5)
which determines a lower bound for $\mathbb{P}_{\mathsf{F81}}\left(y^{(i)}\mid f_{W}(x^{(i)}),T^{(i)}\right)$ in Eq. (3) and, therefore, an upper bound for Eq. (4) itself. We trained PhyloGPN to minimize this bound.
@@ 3.3 Architecture and Training
Figure 2: Architecture of PhyloGPN. A convolutional layer with $N$ input channels, $M$ output channels, a kernel size $K$, and a dilation rate $R$ is noted as having hyperparameters $[N,M,K,R]$. The acronym “RCE” stands for “reverse-complement equivariant.”
@@ 3.4 Embeddings
In certain cases where the relatively small receptive field of PhyloGPN was limiting, we found it beneficial to expand its effective receptive field by pooling embeddings of contiguous positions. For a given integer $\ell\geq 0$, let $\varphi_{-\ell},\dotsc,\varphi_{\ell}\in\mathbb{R}^{d}$ be a set of $d$-dimensional embeddings for the $12\ell+$ 1 positions flanking and including a given position in the center, zero padding whenever necessary, and let $A$ be a fixed $d\times d$ random matrix with entries that are i.i.d. as $\mathcal{N}(0,1).$ We found it effective to apply the transformation
@@ 4.1 Log Likelihood Ratio (LLR) Evaluation
Splice Donor 0.83 0.68 0.67 0.52     0.83
Start Codon 0.77 0.41 0.69 0.43     0.80
Stop Lost 0.81 0.43 0.55 0.48     0.86
Synonymous 0.85 0.55 0.55 0.43     0.87
Upstream of Transcript 0.74 0.40 0.31 0.51     0.84
@@
Predicting deep mutational scanning (DMS) outcomes. We also examined how well LLRs from gLMs can rank the results of DMS assays. These assays evaluate the effects of substitutions on the function of a protein; their results indicate whether substitutions are deleterious. Our data were obtained from various published studies and compiled by Schubach et al. [29]. Our results are summarized in Figure 4 and fully described in Table S1. For 22 out of 25 proteins, PhyloGPN outperforms all baseline models, mostly by a wide margin.
@@ 4.2 Embedding Evaluation
Our results show that PhyloGPN has relatively powerful and broad transfer learning capabilities. Most strikingly, on the Disease VEP task, PhyloGPN obtains an AUROC of 0.96. This is an improvement of 0.19 over the next-best model (Nucleotide Transformer) and of 0.41 over the median. In addition, PhyloGPN achieves state-of-the-art performance on Chromatin Accessibility, Histone Modification, and CpG Methylation tasks, matching or surpassing both gLM baselines and Expert Methods.
@@
However, PhyloGPN does not achieve state-of-the-art performance on certain tasks in BEND. On the Gene Finding task, for instance, PhyloGPN achieves the median result. However, on this task PhyloGPN-X matches Nucleotide Transformer, the previously best-performing gLM. The confusion matrices shown in Figure 5 suggest that a larger context size helps PhyloGPN-X discriminate between intronic sites on the forward and reverse strands. Since these sites are easily distinguished based on their proximity to splice sites, we surmise that having a larger context size makes this information available to the classifier. Interestingly, based on Figure 5, PhyloGPN seems to achieve better performance at discriminating splice donor and splice acceptor sites than PhyloGPN-X or Nucleotide Transformer (Figure S1). We suspect that this is due either to the classifier used by BEND being underpowered or the training set being too limited, rather than because the embeddings provide any less information. Indeed, out of 35,749,767 positions in the training data, only 0.11% are splice sites, whereas 12.45% are in exons, 48.34% are in introns, and 39.10% are in intergenic regions.
@@
Finally, on the Expression VEP task, PhyloGPN only achieves chance-level performance—a seemingly perplexing result given its strong performance on the Disease VEP task. We hypothesized that PhyloGPN has learned to generate similar embeddings for elements that differ by an eQTL because these elements tend to have similar fitness effects. For example, two regulatory elements that bind to the same set of transcription factors may be evolutionary constrained in similar ways despite having different binding affinities. On the other hand, elements that differ by a loss-of-function variant—e.g., a gene and its paralogous pseudogene—are constrained very differently because nonfunctional elements evolve neutrally. To test this hypothesis, we evaluated how well the LLRs generated by PhyloGPN could classify variants in the Expression VEP dataset. On this evaluation, PhyloGPN achieved an AUROC of 0.55. By comparison, Nucleotide Transformer—the best-performing model on the Expression VEP task that we could evaluate in a comparable way—achieved an AUROC of 0.36 on this evaluation. These results imply that, despite its result on the Expression VEP task, PhyloGPN may be in some respects more suitable to perform certain downstream tasks that pertain to predicting variant effects on gene regulation.
On three tasks, GPN-MSA shows relatively worse performance than PhyloGPN. The results on the Gene Finding task are likely caused by GPN-MSA having a relatively tiny 128 bp receptive field. On the other hand, the results on the VEP tasks show that even though GPN-MSA’s LLRs capably predict the labels of variants in BEND’s test sets, the cosine distances between its embeddings of reference and alternate sequences do not. However, recent work has shown that valuable information can still be extracted from GPN-MSA embedding distances across specific latent dimensions [2].
@@ 0.A.3 Evaluation of GPN-MSA on BEND
To obtain embeddings for the BEND tasks, we first used the genome coordinates provided for each sequence to retrieve the corresponding alignments from the multiple sequence alignment (MSA) generated by GPN-MSA. Because the sequences in the BEND tasks exceed GPN-MSA’s context size of 128 bp, we divided each alignment into non-overlapping 128 bp windows, computed embeddings for each window individually, and then concatenated these window embeddings to reconstruct the full-sequence embedding. Note that, since the MSA was constructed using the hg39 human reference genome, we could only evaluate GPN-MSA on all BEND tasks except Chromatin Accessibility, which is based on hg19 coordinates. After obtaining the GPN-MSA embeddings, we followed the default BEND configurations for training and evaluating both the classification tasks and the variant effect prediction tasks.
0.A.4 Supplementary Tables and Figure
@@
A4 0.32 0.50 --0.08-- --0.04--     0.37
ADRB2 0.29 0.30 0.05 0.04     0.43
BRCA1 0.23 --0.09-- --0.18-- --0.10--     0.36
CALM1 0.15 0.00 --0.01-- 0.00     0.15
CYP2C9 (Abundance) 0.26 0.04 --0.05-- 0.03     0.49
CYP2C9 (Activity) 0.24 0.04 --0.03-- 0.01     0.50
DLG4 0.04 0.24 --0.09-- 0.29     0.51
GRB2 0.28 0.24 --0.06-- 0.10     0.31
KCNH2 0.38 --0.07-- 0.19 --0.02--     0.29
MAPK1 0.11 0.08 --0.26-- --0.10--     0.15
MSH2 0.17 0.06 0.03 --0.03--     0.27
NUD15 0.09 0.02 0.09 --0.04--     0.53
TP53 (Null, Etoposide)* 0.38 0.02 0.07 --0.10--     0.39
TP53 (Null, Nutlin)* 0.39 0.02 0.08 --0.08--     0.35
TP53 (WT, Nutlin)* 0.47 --0.04-- --0.02-- --0.11--     0.45
TP53 (Kotler et al. 2018) 0.26 0.01 0.02 --0.06--     0.50
PTEN (Matreyek et al. 2021) 0.17 0.08 0.05 0.00     0.26
PTEN (Mighell et al. 2018) 0.12 0.07 0.02 --0.04--     0.23
SLC6A4 0.24 0.24 --0.06-- --0.00--     0.37
SCN5A 0.23 0.10 --0.15-- 0.02     0.18
SRC 0.20 0.11 --0.02-- --0.01--     0.16
SUMO1 0.13 --0.11-- 0.01 --0.08--     0.17
SNCA 0.19 0.06 0.12 0.00     0.29
TARDBP 0.40 0.11 --0.03-- --0.03--     0.48
TPK1 0.21 0.02 --0.05-- 0.02     0.22
TPMT 0.32 0.05 0.07 --0.04--     0.46
TPOR 0.49 0.15 0.08 --0.03--     0.36
UBE2I 0.13 --0.10-- --0.05-- --0.11--     0.20
VKOR1 (Abundance) 0.25 0.02 0.22 --0.02--     0.39

### v2

@@ 1 Introduction
In recent work, Benegas et al. [1] introduced a gLM named GPN-MSA that achieves cutting-edge performance at this task, even when compared with well-established variant effect predictors. These results are due to a unique aspect of the design of GPN-MSA: it utilizes a multiple sequence alignment (MSA) of diverse vertebrate genomes to the human reference genome as supplementary features, demonstrating that these data are an invaluable resource for modeling human genomes. However, for making a prediction, it requires as input an MSA for the same set of species as in training data, which makes it difficult to apply the model to other species or to regions where alignment to the human genome is poor. This undermines its utility for transfer learning. Furthermore, if genomes that are too similar to the human genome are included in the MSA, the model learns to simply copy from them, resulting in the learned probability distribution not being useful for variant effect prediction. Therefore, most primates were excluded from GPN-MSA’s training data, which is not ideal since they provide useful information about relatively recent evolutionary constraints relevant to humans [15]. Lastly, GPN-MSA relies on existing conservation annotations [32, 27] to filter training data and re-weight the loss function in order to mitigate the influence of poorly conserved regions on the model.
@@ 2.2 Molecular Phylogenetics
The basic data that represent evolutionary relationships between aligned sequences are phylogenetic trees. The leaves of the tree correspond to extant species that are included in the alignment. The interior nodes, on the other hand, correspond to extinct ancestral species. In our context, edges are labeled with values in $0,\infty)$ that represent the evolutionary distances between nodes.
@@ 3.2 Loss Function
$\mathbb{P}_{\mathsf{F81}}(a\mid\theta,b,t)\geq\operatorname{sigmoid}\left(\log t+\sum_{a\in\mathcal{N}}\theta_{a}\right)\pi_{a}\quad\text{for}\ a\neq b,$ (5)
which determines a lower bound for off-diagonal transition probabilities in Eq. (3) and, therefore, an upper bound for the corresponding terms in Eq. (4). We trained PhyloGPN to minimize this bound.
@@ 3.3 Architecture and Training
Figure 2: Architecture of PhyloGPN. A convolutional layer with NN input channels, $M$ output channels, a kernel size $K$, and a dilation rate $R$ is noted as having hyperparameters $[N,M,K,R]$. The acronym “RCE” stands for “reverse-complement equivariant.”
@@ 3.4 Embeddings
In certain cases where the relatively small receptive field of PhyloGPN was limiting, we found it beneficial to expand its effective receptive field by pooling embeddings of contiguous positions. For a given integer $\ell\geq 0$, let $\varphi_{-\ell},\dotsc,\varphi_{\ell}\in\mathbb{R}^{d}$ be a set of $d$-dimensional embeddings for the $2\ell+1$ positions flanking and including a given position in the center, zero padding whenever necessary, and let $A$ be a fixed $d\times d$ random matrix with entries that are i.i.d. as $\mathcal{N}(0,1).$ We found it effective to apply the transformation
@@ 4.1 Log Likelihood Ratio (LLR) Evaluation
Splice Donor 0.82 0.68 0.67 0.52     0.83
Start Codon 0.80 0.41 0.69 0.43     0.80
Stop Lost 0.82 0.43 0.55 0.48     0.86
Synonymous 0.86 0.55 0.55 0.43     0.87
Upstream of Transcript 0.75 0.40 0.31 0.51     0.84
@@
Predicting deep mutational scanning (DMS) outcomes. We also examined how well LLRs from gLMs can rank the results of DMS assays. These assays evaluate the effects of substitutions on the function of a protein; their results indicate whether substitutions are deleterious. Our data were obtained from various published studies and compiled by Schubach et al. [29]. Our results are summarized in Figure 4 and fully described in Table S1. For 24 out of 25 proteins, PhyloGPN outperforms all baseline models, mostly by a wide margin.
@@ 4.2 Embedding Evaluation
Our results show that PhyloGPN has relatively powerful and broad transfer learning capabilities. Most strikingly, on the Disease VEP task, PhyloGPN obtains an AUROC of 0.98. This is an improvement of 0.21 over the next-best model (Nucleotide Transformer) and of 0.43 over the median. In addition, PhyloGPN achieves state-of-the-art performance on Chromatin Accessibility, Histone Modification, and CpG Methylation tasks, matching or surpassing both gLM baselines and Expert Methods.
@@
However, PhyloGPN does not achieve state-of-the-art performance on certain tasks in BEND. On the Gene Finding task, for instance, PhyloGPN performs below the median. However, on this task PhyloGPN-X surpasses Nucleotide Transformer, the previously best-performing gLM. The confusion matrices shown in Figure 5 suggest that a larger context size helps PhyloGPN-X discriminate between intronic sites on the forward and reverse strands. Since these sites are easily distinguished based on their proximity to splice sites, we surmise that having a larger context size makes this information available to the classifier. On the other hand, based on Figure 5, PhyloGPN-X entirely fails to identify splice donor and splice acceptor sites on the forward strand, misclassifying them as exonic sites. PhyloGPN, in contrast, identifies these sites with approximately 90% accuracy. We suspect that this discrepancy is due either to the classifier used by BEND being underpowered or the training set being too limited, rather than because the embeddings provide any less information. Indeed, out of 35,749,767 positions in the training data, only 0.11% are splice sites, whereas 12.45% are in exons, 48.34% are in introns, and 39.10% are in intergenic regions.
@@
Finally, on the Expression VEP task, PhyloGPN performs poorly—a seemingly perplexing result given its strong performance on the Disease VEP task. We hypothesized that PhyloGPN has learned to generate similar embeddings for elements that differ by an eQTL because these elements tend to have similar fitness effects. For example, two regulatory elements that bind to the same set of transcription factors may be evolutionary constrained in similar ways despite having different binding affinities. On the other hand, elements that differ by a loss-of-function variant—e.g., a gene and its paralogous pseudogene—are constrained very differently because nonfunctional elements evolve neutrally. To test this hypothesis, we evaluated how well the LLRs generated by PhyloGPN could classify variants in the Expression VEP dataset. On this evaluation, PhyloGPN achieved an AUROC of 0.53. By comparison, Nucleotide Transformer—the best-performing model on the Expression VEP task that we could evaluate in a comparable way—achieved an AUROC of 0.36 on this evaluation. These results imply that, despite its result on the Expression VEP task, PhyloGPN may be in some respects more suitable to perform certain downstream tasks that pertain to predicting variant effects on gene regulation.
On four tasks, GPN-MSA shows relatively worse performance than PhyloGPN. The results on the Gene Finding task are likely caused by GPN-MSA having a relatively tiny 128 bp receptive field. On the other hand, the results on the VEP tasks show that even though GPN-MSA’s LLRs capably predict the labels of variants in BEND’s test sets, the cosine distances between its embeddings of reference and alternate sequences do not. However, recent work has shown that valuable information can still be extracted from GPN-MSA embedding distances across specific latent dimensions [2].
@@ 0.A.3 Evaluation of GPN-MSA on BEND
To obtain embeddings for the BEND tasks, we first used the genome coordinates provided for each sequence to retrieve the corresponding alignments from the multiple sequence alignment (MSA) generated by GPN‑MSA. Because the sequences in the BEND tasks exceed GPN‑MSA’s context size of 128 bp, we divided each alignment into non-overlapping 128 bp windows, computed embeddings for each window individually, and then concatenated these window embeddings to reconstruct the full-sequence embedding. Note that, since the MSA was constructed using the hg39 human reference genome, we could only evaluate GPN‑MSA on all BEND tasks except Chromatin Accessibility, which is based on hg19 coordinates. After obtaining the GPN‑MSA embeddings, we followed the default BEND configurations for training and evaluating both the classification tasks and the variant effect prediction tasks.
0.A.4 Supplementary Figure and Tables
@@
A4 0.42 0.50 -0.08- -0.04-     0.37
ADRB2 0.35 0.30 0.05 0.04     0.43
BRCA1 0.30 -0.09- -0.18- -0.10-     0.36
CALM1 0.11 0.00 -0.01- 0.00     0.15
CYP2C9 (Abundance) 0.30 0.04 -0.05- 0.03     0.49
CYP2C9 (Activity) 0.24 0.04 -0.03- 0.01     0.50
DLG4 0.47 0.24 -0.09- 0.29     0.51
GRB2 0.30 0.24 -0.06- 0.10     0.31
KCNH2 0.37 -0.07- 0.19 -0.02-     0.29
MAPK1 0.11 0.08 -0.26- -0.10-     0.15
MSH2 0.18 0.06 0.03 -0.03-     0.27
NUD15 0.16 0.02 0.09 -0.04-     0.53
TP53 (Null, Etoposide)* 0.34 0.02 0.07 -0.10-     0.39
TP53 (Null, Nutlin)* 0.34 0.02 0.08 -0.08-     0.35
TP53 (WT, Nutlin)* 0.50 -0.04- -0.02- -0.11-     0.45
TP53 (Kotler et al. 2018) 0.29 0.01 0.02 -0.06-     0.50
PTEN (Matreyek et al. 2021) 0.24 0.08 0.05 0.00     0.26
PTEN (Mighell et al. 2018) 0.15 0.07 0.02 -0.04-     0.23
SLC6A4 0.31 0.24 -0.06- -0.00-     0.37
SCN5A 0.11 0.10 -0.15- 0.02     0.18
SRC 0.15 0.11 -0.02- -0.01-     0.16
SUMO1 0.15 -0.11- 0.01 -0.08-     0.17
SNCA 0.21 0.06 0.12 0.00     0.29
TARDBP 0.36 0.11 -0.03- -0.03-     0.48
TPK1 0.18 0.02 -0.05- 0.02     0.22
TPMT 0.32 0.05 0.07 -0.04-     0.46
TPOR 0.43 0.15 0.08 -0.03-     0.36
UBE2I 0.18 -0.10- -0.05- -0.11-     0.20
VKOR1 (Abundance) 0.24 0.02 0.22 -0.02-     0.39